At January 1, 2021, Café Med leased restaurant equipment from Crescent Corporation under a nine-year lease agreement. The lease agreement specifies annual payments of $21,000 beginning January 1, 2021, the beginning of the lease, and at each December 31 thereafter through 2028. The equipment was acquired recently by Crescent at a cost of $180,000 (its fair value) and was expected to have a useful life of 13 years with no salvage value at the end of its life. (Because the lease term is only 9 years, the asset does have an expected residual value at the end of the lease term of $76,604.) Crescent seeks a 8% return on its lease investments. By this arrangement, the lease is deemed to be an operating lease. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

equired:
1. What will be the effect of the lease on Café Med’s earnings for the first year (ignore taxes)? (Enter decreases with negative sign.)
2. What will be the balances in the balance sheet accounts related to the lease at the end of the first year for Café Med (ignore taxes)?

1. Effect on earnings
2. Lease payable balance (end of year)
Right-of-use asset balance (end of year)
Answer #1

olution 1:
Right to use assets = Present value of lease payments
= $21,000 * cumulative PV factor for annuity due at 8% for 9 periods
= $21,000 * 6.53482 = $141,679

Interest expense for first year = ($141679 - $21,000) * 8% = $11,334
Amortization for the year = $21,000 - $11,334 = $9,666
Effect on earnings for first year = Interest expense + Amortization expense = -$11334 - $9666 = ($21,000)

Solution 2:
Lease payable balance  (End of year) = beginning balance + Interest  expense - Payments
= $141679 + $11334 - $21,000 - $21,000 = $111,014
Right of use asset balance (end of year) = Begininnig balance - Amortization
= $141679 - $9666 = $132,014
Answer #2

Solution 1:
Right to use assets = Present value of lease payments
= $21,000 * cumulative PV factor for annuity due at 8% for 9 periods
= $21,000 * 6.74664 = $141,679

Interest expense for first year = ($141679 - $21,000) * 8% = $11,334
Amortization for the year = $141,769/ 9= $15,742
Effect on earnings for first year = Interest expense + Amortization expense = -$11334 - $15742 = ($27,076)

Solution 2:
Lease payable balance  (End of year) = beginning balance + Interest  expense - Payments
= $141679 + $11334 - $21,000 - $21,000 = $111,014
Right of use asset balance (end of year) = Beginning balance - Amortization
= $141679 - $15742 = $125,937